OpenStudy (anonymous):

giving medals!!! Two cars start at the same place and head in opposite directions. The speed of the first car is 9 less than the speed of the second car. In 1 hour the two cars are 175 miles apart. Find the speed of the second car.

OpenStudy (anonymous):

184

OpenStudy (anonymous):

1 X ( x+ x+9) = 175 2x + 9 = 175 2x = 166 x = 83 mph x+9 = 92 mph ANSWER

OpenStudy (anonymous):

or????84

OpenStudy (anonymous):

good job @alyygirl

OpenStudy (anonymous):

which answer??

OpenStudy (anonymous):

yours

OpenStudy (anonymous):

92 or 84?

OpenStudy (anonymous):

92

OpenStudy (anonymous):

 distance 1 + distance 2 = 175, assuming d=r(t) d1 = (R-9) (1) = R-9 d2= (R)(1)= R Plug in : d1+d2= 175 R-9 + R = 175 2R = 184 R=92 mph

OpenStudy (anonymous):

ok thanks ":)